OpenStudy (anonymous):

please help! use the zero-product property to solve the equations please show your work 1. (x - 9)(x - 8) = 0 2. (7x + 2)(5x - 4) = 0

OpenStudy (anonymous):

1. x-9=0 or x-8=0 so , x=9 or x=8 2. 7x+2=0 or 5x-4=0 so, x=-2/7 or x=4/5

OpenStudy (accessdenied):

The statement of the zero-product property is that, if a product A*B = 0, then either A = 0 or B = 0. This makes sense because anything times zero is equal to zero: A*0 = 0, 0* B = 0 In the question, these are your products: \( \underbrace{(x - 9)}_{one} \underbrace{(x-8)}_{two} = 0\) So we can separate it into this statement: \(x - 9 = 0\) OR \(x - 8 = 0\) Because if either one equals zero, we have: \( 0 * (x - 8) = 0 \) OR \( (x - 9) * 0 = 0 \) Does that make sense?

OpenStudy (anonymous):

okay i got it thanks both of you

OpenStudy (accessdenied):

You're welcome! :)